# XiYan-SQL: A Novel Multi-Generator Framework For Text-to-SQL

### v1

@@ IV-J Study of Generalization
XiYan-SQL is a multi-step Text-to-SQL framework, and it is imperative to discuss its generalization performance across a broader range of out-of-domain datasets, as this is an important indicator of its ability to address SQL generation challenges. In this section, we elaborate on this aspect from two perspectives.
First, from the perspective of the BIRD benchmark, as presented in Table I, XiYan-SQL demonstrates an improvement of 2.3% on the black-box test set compared to the publicly available development set. In particular, this enhancement exceeds that of all previous SOTA methods, which have not surpassed a 2% improvement, with many falling below 1%.
Second, to further illustrate the generalization performance of our fine-tuned model and framework, we present two challenging multi-dialect evaluation sets constructed by our internal data team based on real-world scenarios. These two evaluation sets are utilized to rigorously test the performance of XiYan-SQL on out-of-domain data before the product launch. We present the evaluation results as shown in Table XIII. The PG-Set is a dynamically maintained test set for the PostgreSQL dialect, currently comprising 1,051 question-SQL samples and four large databases. In contrast, the MYSQL-Set test set consists of 1,068 question-SQL samples in the MySQL dialect, along with four databases. These samples, used for evaluation, closely resemble real user business scenarios and encompass multiple dimensions where SQL writing is prone to errors, including date and time calculations (e.g., year-on-year, month-on-month, and trend-related questions), advanced functions, metric statistics, and so on.
Our single fine-tuned model surpasses the baseline (Qwen2.5-Coder-32B) by 5% and 5.6% on two different test sets, respectively, while achieving performance comparable to that of closed-source LLMs, such as GPT and Gemini. The complete XiYan-SQL framework achieves the performance of 50.04% on PG-Set and 58.05% on MYSQL-Set, representing the highest performance, thereby demonstrating robust generalization capabilities. We will also release de-identified versions of these evaluation datasets to contribute to future research in the field. Overall, XiYan-SQL not only focuses on academic research in the domain of SQL generation, but also provides product services to numerous clients in the industry.
TABLE XIII: Comparison of Execution Accuracy performance on the two out-of-domain evaluation sets.
Methods PG-Set(%) MYSQL-Set(%)
GPT-4o 45.93 55.87
Gemini-1.5-pro 45.07 51.33
Qwen2.5-Coder-32B 40.38 46.59
SFT generator (Qwen2.5-Coder-32B) 45.31 52.18
XiYan-SQL (Full pipeline) 50.04 58.05
@@ VI Conclusion
This study presents a novel Text-to-SQL framework, XiYan-SQL, which focuses on generating high-quality and diverse candidate SQL queries, effectively selecting the best among them. Specifically, we introduce a Schema Filter module that employs multi-path retrieval and iterative selection to obtain multiple high-quality schemas. Then, we leverage the proposed multi-task and multi-format fine-tuning strategy to develop a variety of SQL generators, each with distinct advantages, enabling the generation of high-quality and diverse SQL candidates. Finally, we employ a selection model enhanced by a strategy based on model attention to identify the optimal SQL. XiYan-SQL framework achieves SOTA performance on the well-known Text-to-SQL benchmarks, including BIRD and Spider.
There are numerous opportunities for future research on the XiYan-SQL framework. Although the multiple fine-tuned models constructed by this method effectively alleviates the issue of candidate diversity present in traditional fine-tuning approaches, the inherent constraints mean that the diversity it can achieve is still limited. Therefore, we also introduce a generation method based on in-context learning. In the future, we plan to enhance the cognitive capabilities of the fine-tuned models by incorporating additional fine-tuning strategies, thereby further addressing this issue. Moreover, we are currently investigating the integration of a broader range of tasks into an all-in-one SQL model, with the aim of expanding the XiYan-SQL model to encompass more relevant scenarios beyond SQL generation.

### v2

@@ IV-J Study of Generalization
A key indicator of a Text-to-SQL framework’s practical utility is its ability to generalize to diverse, unseen data and environments. We demonstrate the robust generalization of XiYan-SQL from two primary perspectives, both grounded in public, verifiable benchmarks.
First, we examine its in-distribution generalization within the BIRD benchmark. As presented in Table I, XiYan-SQL achieves a 2.3% higher execution accuracy on the black-box test set compared to the public dev set. This improvement margin is notably larger than that of previous state-of-the-art methods, which typically remain below 1%. This result suggests that our framework learns generalizable translation patterns rather than merely overfitting to the development set.
Second, to provide compelling evidence of out-of-domain generalization, we evaluated XiYan-SQL on the BIRD-CRITIC benchmark [14]. BIRD-CRITIC is a novel public benchmark, jointly developed by the BIRD team and Google Cloud, designed to diagnose and resolve complex user challenges in real-world applications. It represents a forward-looking trend in the community, moving beyond simple query translation to address practical, error-prone scenarios. We chose BIRD-CRITIC because it serves as a more rigorous and meaningful litmus test for a model’s practical applicability than older datasets. As verified by the official public leaderboard  [^6] https://bird-critic.github.io/ (as of May 2025), XiYan-SQL achieved a score of 41.0% on this challenging benchmark, reaching SOTA performance. This outstanding result strongly validates that our approach is effective and adaptable in complex, realistic settings. For a detailed description of the experimental setup, the specific adaptations made for the benchmark, and a snapshot of the leaderboard, please refer to Appendix VI.
In summary, the combination of strong performance on the BIRD test set and SOTA results on the public BIRD-CRITIC benchmark confirms that XiYan-SQL is not only a high-performing but also a highly generalizable framework, ready for real-world application challenges.
@@ V Discussion
  • In contrast, API-based approaches that leverage powerful reasoning models (e.g., o4-mini) exhibit superior generalization on certain SQL-reasoning tasks—for example, the “SQL debug” task on the BIRD-CRITIC leaderboard. Our single fine-tuned model, however, demonstrates competitive performance on standard NL2SQL tasks such as BIRD and Spider. To further improve results, we apply the full ensemble pipeline of XiYan-SQL, which enables us to outperform even these strong reasoning models. As detailed in Appendix V, a direct comparison on two complex internal datasets shows that while top-tier API models can surpass our standalone fine-tuned model, our full XiYan-SQL pipeline still achieves the highest accuracy, empirically validating our ensemble framework’s effectiveness.
@@ VI Conclusion
This study presents a novel Text-to-SQL framework, XiYan-SQL, which focuses on generating high-quality and diverse candidate SQL queries, effectively selecting the best among them. Specifically, we introduce a Schema Filter module that employs multi-path retrieval and iterative selection to obtain multiple high-quality schemas. Then, we leverage the proposed multi-task and multi-format fine-tuning strategy to develop a variety of SQL generators, each with distinct advantages, enabling the generation of high-quality and diverse SQL candidates, which we demonstrate is crucial for solving complex queries where a single generation style may fail. Finally, we propose an integrated selection module where a candidate reorganization strategy works in concert with a fine-tuned model to identify the optimal SQL. XiYan-SQL framework achieves SOTA performance on the well-known Text-to-SQL benchmarks, including BIRD and Spider.
Our work highlights the power of an ensemble approach that synergizes the strengths of both specialized fine-tuned models and large-scale ICL models. This combination effectively mitigates the diversity limitations inherent in any single generation method. Looking forward, our framework offers several exciting avenues for future research. We plan to enhance the intrinsic reasoning capabilities of our specialized models, potentially by integrating more sophisticated fine-tuning strategies like self-reflection or constitutional AI principles. Furthermore, we aim to expand the scope of XiYan-SQL beyond pure generation, exploring its potential as a unified “all-in-one” model capable of handling a broader range of SQL-related tasks such as schema linking and query optimization.